OpenStudy (barrelracer17):

Find the equation of the circle whose center and radius are given. center ( 7, -3), radius = 7

rvc (rvc):

\[\rm Center(h,k)~Radius(r):(x-h)^2+(y-k)^2=r^2\]

OpenStudy (michele_laino):

hint: the general equation of a circumference whose center is: \[\Large \left( {{x_0},{y_0}} \right)\] and radius \[\Large r\], is: \[\Large {\left( {x - {x_0}} \right)^2} + {\left( {y - {y_0}} \right)^2} = {r^2}\]

OpenStudy (barrelracer17):

Thank you :)

OpenStudy (michele_laino):

:)